OpenStudy (3psilon):

In a list of 100 numbers , the fraction of numbers greater than 15 is (19/25) and the fraction of numbers less than 15 is (7/50). How many numbers in the list are equal to 15

OpenStudy (3psilon):

@Hero

OpenStudy (3psilon):

Stupid mistakes :P

OpenStudy (kropot72):

Let the quantity of numbers equalling 15 = x \[(100\times \frac{19}{25})+(100\times \frac{7}{50})+x=100\] Now you just need to solve for x.